OpenStudy (anonymous):

PLEASE!!!!!HELP!!!! Find the volume of the solid obtained by rotating the region bounded by the given curves about the specified line. y=x^4, y=1 about y=5

OpenStudy (amistre64):

i like to modify it so that my rotation line is an axis; so lets subtract 5 from all the y parts

OpenStudy (amistre64):

y = x^4 - 5, y=1-5, and about the y=0 line (the xaxis)

OpenStudy (amistre64):